OpenStudy (anonymous):

whats an equation I can use were I Add 7 to both sides. Then, divide both sides by 4.

OpenStudy (anonymous):

really need help

OpenStudy (anonymous):

it is not really clear what the question is asking, but if you start with say \[4x-7=11\] then you would solve by adding 7 to both sides and then dividing both sides by 4

OpenStudy (anonymous):

the 11 of course i just made up, you can use any number there

OpenStudy (anonymous):

would that ecuation equel 11

OpenStudy (anonymous):

i have it equal to 11, but if you solve for \(x\) you do not get \(x=11\) add 7 divide by 4

OpenStudy (amistre64):

im sure that you can add 7 to both sides of any equation, and also divide both sides by 4 an equation remains equal (in balance) as long as you do the same thing to each side.

OpenStudy (anonymous):

all I really want is just an equation

OpenStudy (amistre64):

then any equation will do

OpenStudy (amistre64):

3x+3y-z = 42rx - 8 3x+3y-z + 7 = 42rx - 8 + 7 (3x+3y-z + 7)/4 = (42rx - 8 + 7)/4

OpenStudy (anonymous):

wow I was hoping something more simple

OpenStudy (amistre64):

x = 3 x+7 = 3+7 (x+7)/4 = (3+7)/4 not really sure what you are needing it for tho

OpenStudy (anonymous):

school